# Lifelong Sequential Knowledge Editing without Model Degradation

### v1

@@ Appendix G Norm growth Result
(b) Norm growth of GPT2-XL using EMMET
(c) Norm growth of GPT2-XL using MEMIT
Figure 15: Norm growth of GPT2-XL across different methods
(a) Norm growth of GPT2-XL using MEMIT
(b) Norm growth of GPT2-XL using Norm Constraint
(c) Norm growth of GPT2-XL using ENCORE
Figure 16: Norm growth of GPT2-XL using Norm Constraint and ENCORE
@@
(b) Norm growth of Llama2-7B using EMMET
(c) Norm growth of Llama2-7B using MEMIT
Figure 17: Norm growth of Llama2-7B across different methods
(a) Norm growth of Llama2-7B using MEMIT
(b) Norm growth of Llama2-7B using Norm Constraint
(c) Norm growth of Llama2-7B using ENCORE
Figure 18: Norm growth of Llama2-7B using Norm Constraint and ENCORE
@@
(b) Norm growth of Llama3-8B using EMMET
(c) Norm growth of Llama3-8B using MEMIT
@@
(a) Norm growth of Llama3-8B using MEMIT
(b) Norm growth of Llama3-8B using Norm Constraint
(c) Norm growth of Llama3-8B using ENCORE
Figure 20: Norm growth of Llama3-8B using Norm Constraint and ENCORE
Appendix H Downstream Performance
In this section, we show the downstream performance for each method and also for different algorithms. Note that the reason why x -axis shows the number of edits to be from 0 to 100 is because it shows a number of gradient updates (or batch index) instead of a number of facts edited. Since editing is done in batches of 100, to achieve 10,000 edits we perform a total of 100 batched updates to the models sequentially. Thus, to get the total number of edits finished until a certain point, the x-axis needs to be multiplied by 100, which is the batch size.
Figures 21(a) - 26(a) show the result for the downstream performance for GPT2-XL on both CounterFact and zsRE datasets. Figures 27 - 32 show the result for Llama2-7B on both datasets. Lastly, figures 33 - 38 show the result for Llama3-8B on both datasets.
(a) Downstream Performance for GPT2-XL using AlphaEdit
(b) Downstream Performance for GPT2-XL using AlphaEdit with MPES
Figure 21: Downstream Performance for GPT2-XL using AlphaEdit and MPES with CounterFact dataset
(a) Downstream Performance for GPT2-XL using EMMET
(b) Downstream Performance for GPT2-XL using EMMET with MPES
Figure 22: Downstream Performance for GPT2-XL using EMMET and MPES with CounterFact dataset
(a) Downstream Performance for GPT2-XL using MEMIT
(b) Downstream Performance for GPT2-XL using MEMIT with MPES
(c) Downstream Performance for GPT2-XL using MEMIT with Norm Constraint
(d) Downstream Performance for GPT2-XL using MEMIT with ENCORE
Figure 23: Downstream Performance for GPT2-XL using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
(a) Downstream Performance for GPT2-XL using AlphaEdit
(b) Downstream Performance for GPT2-XL using AlphaEdit with MPES
Figure 24: Downstream Performance for GPT2-XL using AlphaEdit and MPES with zsRE dataset
(a) Downstream Performance for GPT2-XL using EMMET
(b) Downstream Performance for GPT2-XL using EMMET with MPES
Figure 25: Downstream Performance for GPT2-XL using EMMET and MPES with zsRE dataset
(a) Downstream Performance for GPT2-XL using MEMIT
(b) Downstream Performance for GPT2-XL using MEMIT with MPES
(c) Downstream Performance for GPT2-XL using MEMIT with Norm Constraint
(d) Downstream Performance for GPT2-XL using MEMIT with ENCORE
Figure 26: Downstream Performance for GPT2-XL using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
(a) Downstream Performance for Llama2-7B using AlphaEdit
(b) Downstream Performance for Llama2-7B using AlphaEdit with MPES
Figure 27: Downstream Performance for Llama2-7B using AlphaEdit and MPES with CounterFact dataset
(a) Downstream Performance for Llama2-7B using EMMET
(b) Downstream Performance for Llama2-7B using EMMET with MPES
Figure 28: Downstream Performance for Llama2-7B using EMMET and MPES with CounterFact dataset
(a) Downstream Performance for Llama2-7B using MEMIT
(b) Downstream Performance for Llama2-7B using MEMIT with MPES
(c) Downstream Performance for Llama2-7B using MEMIT with Norm Constraint
(d) Downstream Performance for Llama2-7B using MEMIT with ENCORE
Figure 29: Downstream Performance for Llama2-7B using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
(a) Downstream Performance for Llama2-7B using AlphaEdit
(b) Downstream Performance for Llama2-7B using AlphaEdit with MPES
Figure 30: Downstream Performance for Llama2-7B using AlphaEdit and MPES with zsRE dataset
(a) Downstream Performance for Llama2-7B using EMMET
(b) Downstream Performance for Llama2-7B using EMMET with MPES
Figure 31: Downstream Performance for Llama2-7B using EMMET and MPES with zsRE dataset
(a) Downstream Performance for Llama2-7B using MEMIT
(b) Downstream Performance for Llama2-7B using MEMIT with MPES
(c) Downstream Performance for Llama2-7B using MEMIT with Norm Constraint
(d) Downstream Performance for Llama2-7B using MEMIT with ENCORE
Figure 32: Downstream Performance for Llama2-7B using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
(a) Downstream Performance for Llama3-8B using AlphaEdit
(b) Downstream Performance for Llama3-8B using AlphaEdit with MPES
Figure 33: Downstream Performance for Llama3-8B using AlphaEdit and MPES with CounterFact dataset
(a) Downstream Performance for Llama3-8B using EMMET
(b) Downstream Performance for Llama3-8B using EMMET with MPES
Figure 34: Downstream Performance for Llama3-8B using EMMET and MPES with CounterFact dataset
(a) Downstream Performance for Llama3-8B using MEMIT
(b) Downstream Performance for Llama3-8B using MEMIT with MPES
(c) Downstream Performance for Llama3-8B using MEMIT with Norm Constraint
(d) Downstream Performance for Llama3-8B using MEMIT with ENCORE
Figure 35: Downstream Performance for Llama3-8B using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
(a) Downstream Performance for Llama3-8B using AlphaEdit
(b) Downstream Performance for Llama3-8B using AlphaEdit with MPES
Figure 36: Downstream Performance for Llama3-8B using AlphaEdit and MPES with zsRE dataset
(a) Downstream Performance for Llama3-8B using EMMET
(b) Downstream Performance for Llama3-8B using EMMET with MPES
Figure 37: Downstream Performance for Llama3-8B using EMMET and MPES with zsRE dataset
(a) Downstream Performance for Llama3-8B using MEMIT
(b) Downstream Performance for Llama3-8B using MEMIT with MPES
(c) Downstream Performance for Llama3-8B using MEMIT with Norm Constraint
(d) Downstream Performance for Llama3-8B using MEMIT with ENCORE
Figure 38: Downstream Performance for Llama3-8B using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
@@ Appendix I Editing Performance on zsRE dataset
Model Method Edit Score Paraphrase Score Neighborhood Score
GPT2-XL EMMET 39.71 27.17 9.78
EMMET+MPES 53.55 39.42 16.81
AlphaEdit 42.1 33.61 14.61
AlphaEdit + MPES 54.99 43.18 18.4
MEMIT 74.6 61.77 22.4
MEMIT + MPES 75.09 61.58 23.37
Norm-Constraint 74.51 61.9 23.39
ENCORE 74.46 61.79 23.41
@@
Model Method Edit Score Paraphrase Score Neighborhood Score
Llama2-7B EMMET 75.42 69.69 33.89
EMMET+MPES 84.07 76.9 41.88
AlphaEdit 83.77 77.12 41.96
AlphaEdit + MPES 83.8 77.64 41.97
MEMIT 79.49 74.29 41.8
MEMIT + MPES 83.01 77.45 44.64
Norm-Constraint 88.73 84.05 47.98
ENCORE 89.1 84.28 48.51
@@
Model Method Edit Score Paraphrase Score Neighborhood Score
Llama3-8B EMMET 96.97 90.96 45.26
EMMET+MPES 96.75 91.31 46.44
AlphaEdit 89.27 82.19 45.23
AlphaEdit + MPES 93.54 85.93 47.32
MEMIT 96.45 90.3 48.91
MEMIT + MPES 96.85 90.76 47.34
Norm-Constraint 90.4 84.58 49.09
ENCORE 93.15 86.19 49.81
Appendix J Evaluation of Downstream Performance
@@
In this paper, we assess model degradation by measuring downstream performance at regular intervals of edits. Our evaluation suite is wide-ranging and consists of the following 6 tasks – sentiment analysis (SST2) (Socher et al., 2013), paraphrase detection (MRPC) (Dolan & Brockett, 2005), natural language inference (NLI, RTE) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009), linguistic acceptability classification (CoLA) (Warstadt et al., 2019), and massive multitask language understanding (MMLU) (Hendrycks et al., 2020).
@@
For each task, we created a subset of 100 examples balanced across all multiple-choice options. The models were evaluated on the tasks above, and the accuracy score was measured every 1000 edits. In order to improve the models’ initial performance and achieve meaningful signals, we provided few-shot examples. The few-shot prompt templates used for each task are shown in Figures 39-44.

### v2

@@ Appendix H Norm growth Result
(b) Norm growth of GPT2-XL using MEMIT
Figure 13: Norm growth of GPT2-XL using AlphaEdit and MEMIT
(a) Norm growth of GPT2-XL using NC
(b) Norm growth of GPT2-XL using MPES + NC
Figure 14: Norm growth of GPT2-XL using NC and MPES + NC
@@
(b) Norm growth of Llama2-7B using MEMIT
Figure 15: Norm growth of Llama2-7B using AlphaEdit and MEMIT
(a) Norm growth of Llama2-7B using NC
(b) Norm growth of Llama2-7B using MPES + NC
Figure 16: Norm growth of Llama2-7B using NC and MPES + NC
@@
(b) Norm growth of Llama3-8B using MEMIT
@@
(a) Norm growth of Llama3-8B using NC
(b) Norm growth of Llama3-8B using MPES + NC
Figure 18: Norm growth of Llama3-8B using NC and MPES + NC
@@ Appendix I Editing Performance on zsRE dataset
Model Method Edit Score Paraphrase Score Neighborhood Score GPT2-XL AlphaEdit 42.1 33.61 14.61 AlphaEdit + MPES 54.99 43.18 18.40 MEMIT 74.60 61.77 22.40 MEMIT + MPES 75.09 61.58 23.37 MEMIT + NC 74.51 61.90 23.39 MEMIT + MPES + NC 74.46 61.79 23.41
@@
Model Method Edit Score Paraphrase Score Neighborhood Score Llama2-7B AlphaEdit 83.77 77.12 41.96 AlphaEdit + MPES 83.80 77.64 41.97 MEMIT 79.49 74.29 41.80 MEMIT + MPES 83.01 77.45 44.64 MEMIT + NC 88.73 84.05 47.98 MEMIT + MPES + NC 89.10 84.28 48.51
@@
Model Method Edit Score Paraphrase Score Neighborhood Score Llama3-8B AlphaEdit 89.27 82.19 45.23 AlphaEdit + MPES 93.54 85.93 47.32 MEMIT 96.45 90.30 48.91 MEMIT + MPES 96.85 90.76 47.34 MEMIT + NC 90.40 84.58 49.09 MEMIT + MPES + NC 93.15 86.19 49.81
@@
Appendix J Downstream Performance
Figures 19(a) - 20(b) show the result for the downstream performance for GPT2-XL on both CounterFact and zsRE datasets. Figures 21(a) - 22(b) show the result for Llama2-7B on both datasets. Lastly, figures 23(a) - 24(b) show the result for Llama3-8B on both datasets.
(a) AlphaEdit and MPES
(b) MEMIT, MPES, NC, MPES+NC
Figure 19: Downstream Performance for GPT2-XL using different editing methods with CounterFact dataset
(a) AlphaEdit and MPES
(b) MEMIT, MPES, NC, MPES+NC
Figure 20: Downstream Performance for GPT2-XL using different editing methods with zsRE dataset
(a) AlphaEdit and MPES
(b) MEMIT, MPES, NC, MPES+NC
Figure 21: Downstream Performance for Llama2-7B using different editing methods with CounterFact dataset
(a) AlphaEdit and MPES
(b) MEMIT, MPES, NC, MPES+NC
Figure 22: Downstream Performance for Llama2-7B using different editing methods with zsRE dataset
(a) AlphaEdit and MPES
(b) MEMIT, MPES, NC, MPES+NC
Figure 23: Downstream Performance for Llama3-8B using different editing methods with CounterFact dataset
(a) AlphaEdit and MPES
(b) MEMIT, MPES, NC, MPES+NC
Figure 24: Downstream Performance for Llama3-8B using different editing methods with zsRE dataset
Appendix K Norm Decrease From MPES
In this section we want to highlight one observation that we have is that while MPES does reduce the norm of the edited layer shown in Figure 25 by quite a lot compared to just MEMIT alone in Figure 1(a). But we still note that the norm growth is still 3 times higher which is the reason why we think that explicit norm control is still needed.
Figure 25: The figure shows the Norm-growth as function of number edits in MEMIT Llama3-8B with MPES.
Appendix L Unbound growth of norm in AlphaEdit
Here we provide the result if we remove the norm constraint from the objective function of AlphaEdit. As we can see, once we remove the norm constraint, the norm growth of AlphaEdit just becomes unbound. Note that the reason why the x-axis only goes up to 300 edits is that after that the norm value becomes overflow in Python and it therefore it cannot be plot.
(a) AlphaEdit with Llama2-7B
(b) AlphaEdit with Llama3-8B
Figure 26: The figures show the unbound Norm-growth as function of number edits in AlphaEdit
Appendix M Batch Size Effect
For this section we show the result for different batch size we see that our method MPES + NC is still effective for all batch size and for all model
Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B MEMIT 91.97 81.73 50.32 77.67 65.58 49.56 58.00 66.22 50.59 73.19 70.45 50.15 503.98 522.72 281.88 PRUNE 92.76 51.80 48.50 78.43 50.15 49.21 57.25 50.17 51.15 73.17 50.70 49.60 505.02 247.88 297.80 RECT 51.01 51.85 48.94 48.92 50.47 49.98 52.57 49.34 50.14 50.79 50.53 49.68 539.58 323.46 192.70 AlphaEdit 91.24 56.92 57.80 73.71 51.38 56.55 56.39 53.73 49.70 70.99 53.92 54.44 586.77 494.64 430.37 MEMIT + MPES + NC 92.57 90.79 89.71 78.19 79.76 80.48 60.36 59.70 57.63 74.70 74.44 73.31 510.81 555.18 539.91
Table 15: Editing performance of our approach when compared to baseline MEMIT, AlphaEdit and MEMIT regularization method such as PRUNE and RECT using batchsize 10.
(a) Downstream Performance for GPT2-XL with batch size 10
(b) Downstream Performance for Llama2-7B with batch size 10
(c) Downstream Performance for Llama3-8B with batch size 10
Figure 27: Different Downstream Performance for different models with batch size 10
Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B MEMIT 94.04 81.04 49.68 79.91 64.67 49.29 57.90 60.95 51.31 74.22 67.86 50.08 517.37 442.59 373.48 PRUNE 61.05 70.80 49.38 58.05 62.11 49.63 50.00 51.86 51.09 55.96 60.60 50.02 579.69 280.83 340.22 RECT 51.40 82.42 63.17 49.83 66.84 56.92 52.17 67.39 52.89 51.12 71.54 57.36 409.42 549.35 588.39 AlphaEdit 88.58 61.10 72.67 70.33 55.86 63.44 56.04 53.75 52.90 69.20 56.74 61.95 580.27 540.92 465.81 MEMIT + MPES + NC 93.35 92.57 88.77 78.66 82.64 78.19 59.84 60.43 60.07 74.75 76.04 73.71 523.49 560.16 523.61
Table 16: Editing performance of our approach when compared to baseline MEMIT, AlphaEdit and MEMIT regularization method such as PRUNE and RECT using batchsize 100.
(a) Downstream Performance for GPT2-XL with batch size 100
(b) Downstream Performance for Llama2-7B with batch size 100
(c) Downstream Performance for Llama3-8B with batch size 100
Figure 28: Different Downstream Performance for different models with batch size 100
Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B MEMIT 93.55 93.94 74.45 81.31 76.83 61.25 59.66 67.64 57.33 75.47 78.03 63.56 558.93 577.69 457.97 PRUNE 92.42 88.58 62.83 81.27 70.43 54.11 58.79 64.64 52.29 74.75 73.25 56.05 548.27 527.05 461.71 RECT 87.49 87.35 57.36 71.94 60.86 52.31 64.9 73.67 64.34 73.64 72.37 57.59 603.91 566.67 215.44 AlphaEdit 92.61 96.11 91.29 76.78 87.27 76.88 56.09 61.63 68.95 72.03 78.76 78.00 587.19 588.09 593.35 MEMIT + MPES + NC 92.57 93.75 84.29 80.01 80.45 70.51 61.26 65.71 69.48 75.71 78.30 74.19 566.94 593.12 598.86
Table 17: Editing performance of our approach when compared to baseline MEMIT, AlphaEdit and MEMIT regularization method such as PRUNE and RECT using batchsize 1000.
(a) Downstream Performance for GPT2-XL with batch size 1000
(b) Downstream Performance for Llama2-7B with batch size 1000
(c) Downstream Performance for Llama3-8B with batch size 1000
Figure 29: Different Downstream Performance for different models with batch size 1000
Appendix N Evaluation of Downstream Performance
@@
In this paper, we assess model degradation by measuring downstream performance at regular intervals of edits. Our evaluation suite is wide-ranging and consists of the following 6 tasks – sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI, RTE) Dagan et al. (2005); Haim et al. (2006); Giampiccolo et al. (2007); Bentivogli et al. (2009), linguistic acceptability classification (CoLA) Warstadt et al. (2019), and massive multitask language understanding (MMLU) Hendrycks et al. (2020).
@@
For each task, we created a subset of 100 examples balanced across all multiple-choice options. The models were evaluated on the tasks above, and the accuracy score was measured every 1000 edits. In order to improve the models’ initial performance and achieve meaningful signals, we provided few-shot examples. The few-shot prompt templates used for each task are shown in Figures 30-35.